Differentiate the function or find 𝑑𝑦 𝑑π‘₯ and simplify the result.

1. 𝑓(π‘₯) = ln(4 + 5π‘₯)

2. 𝐹(𝑦) = ln(sin 5𝑦)

3. 𝑦 = ln|π‘₯3 + 1|

4. 𝑦 = π‘₯5(π‘₯+2)/π‘₯βˆ’3 (by logarithmic differentiation)

5. 𝑦 = 𝑒5x

6. 𝑦 = 𝑒x sin 𝑒x

7. 𝑓(π‘₯) = 35x

8. 𝑓(π‘₯) = 4sin2x

9. 𝑓(π‘₯) = sinh π‘₯2

10. 𝑓(π‘₯) = coth 1/x


1
Expert's answer
2021-02-24T06:02:12-0500
  1. "\bold{f(x) = ln(4+5x)}"

Let


"u = 4+5x\\"

Then


"f(x) = ln(u)"

By chain rule,


"\dfrac{d}{dx}[f(x)] = \dfrac{d}{du}[f(x)] \times \dfrac{d}{dx}[u]"

"\dfrac{d}{du}[f(x)] = \dfrac{1}{u} = \dfrac{1}{4+5x}""\dfrac{d}{dx}[u] = 5""\therefore \bold{\dfrac{d}{dx}[f(x)]} = \dfrac{1}{4+5x} \times 5 = \bold{\dfrac{5}{4+5x}}"

"\bold{2. \quad F(y) = \ln(\sin(5y))}"


"\dfrac{d}{dy}[\ln(\sin(5y))]\\
= \dfrac{1}{\sin(5y)} \times \dfrac{d}{dy}[5y]\\
= \dfrac{\cos(5y) \times \frac{d}{dy}[5y]}{\sin(5y)}\\
= \dfrac{\cos(5y) \cdot 5 \cdot \frac{d}{dy}[5]}{\sin(5y)}\\
= \dfrac{5\cos(5y) \cdot 1}{\sin(5y)}\\
\bold{F'(y) = \dfrac{5\cos(5y)}{\sin(5y)}}"

"\bold{3. \quad 𝑦 = \ln|𝑥^3 + 1|}"


"y' = \dfrac{1}{|x^3+1|} \cdot \dfrac{d}{dx}[|x^3+1|]\\
= \dfrac{\frac{x^3+1}{|x^3+1|} \cdot \frac{d}{dx}[x^3+1]}{|x^3+1|}\\
= \dfrac{\frac{d}{dx}[x^3]+\frac{d}{dx}[1]}{x^3+1}\\
= \dfrac{3x^2+0}{x^3+1}\\
\bold{y'=\dfrac{3x^2}{x^3+1}}"